# Large Language Models as Span Annotators

### v2

@@ 4.1 Task: Evaluation of Data-to-text Generation
We gather span annotations for $\mathcal{D}_{\texttt{test}}$ with crowdworkers from Prolific.com. We apply best practices for gathering human annotations, including an iterative process of refining annotation guidelines and preselecting best-performing annotators using a qualification task Tseng et al. (2020); Iskender et al. (2020); Huang et al. (2023); Zhang et al. (2023). Of the 230 annotators that participated in the qualification task, we selected 50 best annotators to annotate our data. We provide more details on the process of collecting human annotations in Section B.2.
4.2 Task: Identifying Errors in Machine Translation
@@
For MT-Eval, we use the dataset of system outputs from the WMT 2024 general shared task (Kocmi et al., 2024b). The system outputs were annotated with the Error Span Annotation (ESA) protocol (Kocmi et al., 2024c) by professional translators. We focus on character-level span annotations of major and minor translation errors provided in the dataset. The inputs $X$ in MT-Eval are the source texts used to generate the translation $Y$ [^8] Note unlike for the other tasks, the annotations in MT-Eval cannot overlap and need not to be aligned with word boundaries.
@@
We use data for the three textual domains present in the WMT 2024 shared task: news, literary, and social. We focus on data translated from English into other languages, including Chinese, Czech, German, Hindi, Icelandic, Japanese, Russian, Spanish, Ukrainian.
@@
The original dataset has nearly 50k model outputs, making it too costly to run the full evaluation campaign. Therefore, we used all available system outputs for 10 sampled inputs for each language, giving us 2,854 outputs that we use for the experiments.
4.3 Task: Propaganda Technique Detection
@@
For the Propaganda task, we use the dataset of Da San Martino et al. (2019). The dataset contains news collected from mostly propagandistic on-line sources. The token-level annotations in the dataset created by expert annotators cover 18 categories of logical fallacies and persuasion techniques. We use the test split for our experiments. Note that there are no inputs $X$ for the task, as all the annotated categories are intrinsic to the evaluated text $Y$.
4.4 Mitigating Data Contamination
We took preventive measures to ensure that our results are not affected by data contamination Balloccu et al. (2024); Dong et al. (2024); Jiang et al. (2024b):
  • For D2T-Eval, we follow the approach of Kasner & Dušek (2024) to download a new set of structured data from publicly available sources and generate new outputs.
  • For MT-Eval, we use the WMT 2024 dataset that was released at the end of 2024, which is after the knowledge cut-off date of the LLMs we are using.
  • For Propaganda, the dataset is publicly available since 2019, but due to its format cannot be easily memorized. [^9] The model would need to separately match the annotations for numerical token ids with the text.
@@ 4.5 Collecting LLM annotations
For our span annotator LLMs, we use a mix of open and proprietary state-of-the-art models: (1) instruction-tuned models Llama 3.3 70B Grattafiori et al. (2024), GPT-4o Hurst et al. (2024), Claude 3.7 Sonnet Anthropic (2025), and (2) reasoning models: DeepSeek-R1 70B DeepSeek-AI et al. (2025), o3-mini OpenAI (2025) and Gemini 2.0 Flash Thinking Deepmind (2025). We run the open models (Llama 3.3, DeepSeek-R1) locally in 4-bit quantization using ollama.
@@
We define several prompt variants for our experiments. $\mathcal{P}_{\texttt{base}}$ is the base prompt that includes the guidelines $\mathcal{G}$ as given to human annotators. By extending $\mathcal{P}_{\texttt{base}}$, we implement few-shot ($\mathcal{P}_{\texttt{5shot}}$) and chain-of-thought ($\mathcal{P}_{\texttt{cot}}$) prompts. We also ablate $\mathcal{P}_{\texttt{base}}$ by removing extended guidelines ($\mathcal{P}_{\texttt{noguide}}$) and not asking for explanations ($\mathcal{P}_{\texttt{noreason}}$). The full prompts can be found in Appendix D.
Annotation Tool
We use factgenie Kasner et al. (2024), a tool that supports collecting span annotations from humans via a web interface and from LLMs via API calls.
@@ 5 Results
In this section, we describe the results of our experiments on span annotation with LLMs. We first investigate the effect of prompting techniques in Section 5.1. Next, we evaluate the LLM annotations using automatic metrics (Section 5.2) and manually analyze the errors in the model outputs (Section 5.3).
@@ 5.1 Prompting Techniques
We perform a small-scale study on $\mathcal{D}_{\texttt{dev}}$ using open models (Llama 3.3 and DeepSeek-R1) to study the differences between prompting techniques. The results are shown in Table 2.
@@
Prompt $F_{1}$ $\gamma$a/o $F_{1}$ $\gamma$a/o
@@
Including detailed guidelines seems beneficial: omitting the guidelines ($\mathcal{P}_{\texttt{noguide}}$) lowers the performance of both models. In contrast, not letting the model explain the annotation ($\mathcal{P}_{\texttt{noreason}}$) does not have a substantial effect. For Llama 3.3, the chain-of-thought (CoT) prompting ($\mathcal{P}_{\texttt{cot}}$) makes it produce less annotations per example than the base variant (0.8 vs. 2.4), leading to lower F1 and $\gamma$ scores. [^10] The model with $\mathcal{P}_{\texttt{cot}}$ tends to “overthink” the annotations, deciding not to annotate cases of errors against which it can find some arguments. We provide an example of this behavior in Table 27. Few-shot prompting ($\mathcal{P}_{\texttt{5shot}}$) brings ambivalent results, increasing Llama 3.3 scores but doing the opposite for DeepSeek-R1.111111This observation is aligned with DeepSeek-AI et al. (2025), who note that few-shot prompting consistently degrades the performance of DeepSeek-R1. Given these considerations, we decided to use $\mathcal{P}_{\texttt{base}}$ for our main experiments.
@@
Figure 2: Comparison of LLMs using $\mathcal{P}_{\texttt{base}}$ and human annotators. The stripe pattern denotes proprietary models, darker shades denote reasoning models. For $F_{1}$ score, the grey bar denotes its soft variant. The dashed horizontal lines for D2T-Eval denote agreement between our human annotators. Detailed results are included in Tables 8, 16 and 20.
Next, we compute IAA between LLM and human annotations using the metrics described in Section 3.3. The overall results for all tasks are given in Figure 2. We also provide detailed results for individual tasks in Appendix E.
Figure 3: Confusion matrix for D2T-Eval (Contradictory, Not checkable, Misleading, Incoherent, Repetitive, Other, cf. Table 5).
Model football gsmarena openweath.
$F_{1}$ $\gamma$ $F_{1}$ $\gamma$ $F_{1}$ $\gamma$
Llama 3.3 0.29 0.20 0.11 0.08 0.12 0.05
GPT-4o 0.30 0.19 0.09 0.06 0.18 0.12
Claude 3.7 0.38 0.25 0.16 0.14 0.27 0.28
DeepS. R1 0.34 0.28 0.12 0.16 0.18 0.12
o3-mini 0.42 0.33 0.15 0.19 0.29 0.31
Gem. 2-FT 0.38 0.31 0.16 0.16 0.22 0.24
Table 3: Performance of models by domain (football, gsmarena, openweather) in D2T-Eval: $F_{1}$ score (hard), $\gamma$ score. The best results are bolded.
@@
DeepSeek-R1 generally outperforms the same-sized Llama 3.3. The superiority of DeepSeek-R1 is most pronounced on D2T-Eval ($F_{1}$-score of 0.23 vs. 0.18, $\gamma$ score of 0.19 vs. 0.11). The same observation applies to OpenAI models, where o3-mini outperforms GPT-4o despite o3-mini’s price per token being approximately 2x lower. A notable exception to this trend is the Claude 3.7 Sonnet, which is mostly on par with o3-mini and excels on MT-Eval. However, its price per token is approximately 3x higher (cf. Table 4 for model costs).
@@
For D2T-Eval, we can compare model results with an average IAA on examples annotated by two human annotators (cf. Section B.2). Here, o3-mini, Claude 3.7 and Gemini 2.0 mostly reach or surpass human agreement. For Propaganda, the upper bound of IAA is the result of Da San Martino et al. (2019), who report $\gamma=0.31$ for annotators before consolidation. This score is substantially higher (the best LLM score being $\gamma=0.16$ for the Claude 3.7 Sonnet). However, this task used expert annotators and has the largest number of categories. The latter manifests itself in the large difference between the soft and hard $F_{1}$ scores. [^12] The WMT24 dataset for MT-Eval contains examples annotated with a pair of annotators, but these examples form only a small fraction and exhibit high variance between language pairs. Therefore, we decided to omit the comparison with human IAA in MT-Eval.
@@
Confusion matrices (cf. Figure 3 and Section E.8) suggest that the models tend mainly to confuse related categories such as Contradictory, Not checkable and Misleading for D2T-Eval, which may be related to ambiguity or subjective understanding of category definitions. The models also use a less diverse distribution of categories, e.g., resorting to the Contradictory category more often than human annotators.
LLMs are more efficient than human annotators.
For D2T-Eval, crowdsourced annotation for 1k outputs costs approximately \$500, while annotating the same amount of outputs with the high-performance model o3-mini LLM costs \$3.60 (cf. Table 4). The crowdworkers take 129.1 seconds per output on average, which is better than DeepSeek-R1 on our local infrastructure, but order of magnitude slower than API-based models. [^13] Also note that unlike LLMs, the crowdworkers are not required to produce explanations for the annotations. This suggests that LLMs can be an effective alternative to human annotators both in terms of costs and time.
Metric crowdw. Llama 3.3 DeepSeek-R1 Claude 3.7 Sonnet o3-mini
Cost per 1k outputs (\$) 500 - - 10.5 3.6
Time per output (s) 129.1 21.6 227.5 9.0 21.8
Table 4: Estimate of costs and time requirements on D2T-Eval: crowdworkers on Prolific, open models (Llama 3.3, DeepSeek-R1), and proprietary models (Claude 3.7 Sonnet, o3-mini). Note that we omit the infrastructure and energy costs for the models deployed locally.
@@ 5.3 Manual Analysis of LLM Annotations
We manually analyzed the quality of LLM annotations on 216 samples from D2T-Eval and Propaganda. For each model, we sampled three annotations per category in D2T-Eval and one annotation per category in Propaganda. We show the results in Figure 4.
In total, we marked 49.5% of LLM-generated annotations and 50.5% of reasons as correct (with 9.2% of annotations and 12.5% of reasons additionally marked as partially correct). Reasoning models perform better, with 56.4% of their annotations and 58.3% marked as completely correct. The most accurate annotations on D2T-Eval were those made by Gemini 2.0 and DeepSeek-R1. o3-mini performed well on both tasks, although Propaganda proved to be challenging for all models.
@@
Qualitatively, we find that the models often select wrong error categories despite identifying real issues (e.g. labeling contradictory statements as "incoherent"). Models also tend to be overly attentive, flagging noise in the data (quotes or non-article content) as errors, or marking slight numerical variations (such as rounded values) as misleading. All of these cases could be tackled by more descriptive guidelines or additional examples. However, in some cases, the models also misread or misinterpreted the data (e.g., claiming wind speed measurements do not exist when they do), which hints at deeper issues with understanding the data. Incorrect explanations vary from incomplete explanations (addressing only part of a multi-issue span), irrelevant explanations (e.g., counting omissions as sources of errors) to incorrectly flagging subjective statements (e.g., “long-lasting usage”) as factual errors. Occasionally, the model admits that it marked a correct span as an error, such as in “The description of the game’s duration aligns with the data, providing coherent information”. [^14] This typically happens to GPT-4o, even though OpenAI API ensures JSON key ordering so the explanation should have been generated before the annotation (cf. Section 3.2).
@@
Interestingly, the LLM annotations that were marked as correct have only 24% hard character-level overlap (51% soft) with human annotations. This fact led us to analyze the quality of human annotations in D2T-Eval. On a sample of 108 human annotations, we marked 45.3% of the annotations and 54.6% of the explanations correct, which is comparable to the LLM annotations. This suggests that the task is hard even for human annotators, and the quality of annotations from crowdworkers varies, even if they are preselected using a qualification task.
Figure 4: Results of our manual analysis. We analyzed 18 annotations and their explanations for each model and task (216 annotations in total). The color bars show annotations that we classified as correct, partially correct, assigned wrong category, incorrect, and undecidable cases. Detailed results are provided in Tables 22 and 23 in Appendix E.
@@ 6 Discussion
The IAA between LLMs and human annotators is generally moderate. However, we also need to take into account a realistic upper bound: the average IAA between human annotators themselves. In that case, the strongest LLMs have already reached the level of skilled crowdworkers in D2T-Eval. As we also showed, not all human annotations are “gold”, as even qualified crowdworkers make similar amounts of errors as LLMs. If we also factor in financial aspects, LLMs provide better flexibility, shorter response times, and lower costs. In light of this, we argue that the advantages of automating span annotation with LLMs outweigh its issues. LLM-based span annotation will also benefit from future increases in LLM capabilities.
Recommendations
To use an LLM as a span annotator, we strongly recommend providing detailed guidelines describing how to handle all edge cases. In contrast, we note that providing specific examples (i.e., a few-shot setup) does not bring consistent improvements, possibly due to the length and complexity of examples for this task. When using LLMs in a new setup, we recommend validating model’s annotations against examples hand-annotated by experts on a sample of the data. Regarding model choice, we recommend choosing a reasoning model as opposed to an equivalent model without reasoning, since the reasoning models provide more reliable annotations at the cost of slightly higher response times and token count.

### v3

@@ 4.1.1 D2T-Eval: Evaluation of Data-to-text Generation
To gather annotations for our dataset, we use crowdworkers from Prolific.com. We apply best practices for gathering human annotations, including an iterative process to refine annotation guidelines and preselecting the best-performing annotators using a qualification task tseng2020best; iskender2020best; huang2023incorporating; zhang-etal-2023-needle. Our process of collecting annotations proceeded in two stages, following the setup of zhang-etal-2023-needle: (1) a qualification task for pre-selecting skilled annotators, and (2) the main task for collecting the annotations. See Appendix B.2 for more details on collecting annotations.
For quality checks, we collect additional internal gold annotation (by the authors) for subsets of the data: $\mathcal{D}_{\texttt{dev}}$ for selecting the best prompt and $\mathcal{D}_{\texttt{iaa}}$ for validating the performance of human annotators (cf. Appendix B.2). Here is a complete overview of our data splits for D2T-Eval:
  • $\mathcal{D}_{\texttt{test}}$ (1200 outputs) – for LLM evaluation, annotated with crowdworkers,
  • $\mathcal{D}_{\texttt{dev}}$ (84 outputs) – for the study of prompt variants, annotated internally,
  • $\mathcal{D}_{\texttt{iaa}}$ (12 outputs) – control for human crowdworkers, annotated internally.
4.1.2 MT-Eval: Identifying Errors in Machine Translation
@@
For MT-Eval, we use the dataset of system outputs from the WMT 2024 general shared task (kocmi2024findings). The system outputs were annotated with the Error Span Annotation (ESA) protocol (kocmi2024error) by professional translators.
The inputs $X$ for MT-Eval are the texts in the source language used to produce the translation $Y$. We follow the WMT 2024 annotation guidelines, focusing on character-level span annotations of Major and Minor translation errors (see Table˜6 for their definitions). Note that unlike for the other tasks, the annotations in MT-Eval cannot overlap and need not be aligned with word boundaries.
@@
We select the three textual domains present in the WMT 2024 shared task: news, literary, and social; using the data translated from English into other languages: Chinese, Czech, German, Hindi, Icelandic, Japanese, Russian, Spanish, and Ukrainian.
@@
The original dataset has nearly 50k model outputs, making it too extensive for our evaluation campaign. Therefore, we used a balanced sample of all available system outputs for 10 sampled inputs for each pair of languages, making up 2,854 examples in total.
4.1.3 Propaganda: Propaganda Technique Detection
@@
For the Propaganda task, we use the dataset of dasanmartino2019finegrained containing news collected mostly from on-line propagandistic sources. The token-level annotations in the dataset created by expert annotators cover 18 categories of logical fallacies and persuasion techniques. We use the test split for our experiments. Inputs $X$ are empty for this task, as all annotated categories are intrinsic to the evaluated text $Y$.
@@ 4.2 Collecting LLM annotations
For collecting span annotations with LLMs, we use a mixture of open and proprietary state-of-the-art models:
  • instruction-tuned models: Llama 3.3 70B grattafiori2024llama, GPT-4o hurst2024gpt, and Claude 3.7 Sonnet anthropic2025claude,
  • reasoning models: [^4] By reasoning models we understand the models that use extra inference time to generate a thinking trace before providing the answer marjanovic2025deepseek. DeepSeek-R1 distilled 70B version guo2025deepseek, o3-mini openai2025o3mini, and Gemini 2.0 Flash Thinking google2025geminithinking.
See Appendix˜A for details on our experimental setup.
@@
We define several prompt variants for our experiments. $\mathcal{P}_{\texttt{base}}$ is the base prompt that includes the guidelines $\mathcal{G}$ as given to human annotators and asks the model to explain its annotation. By extending $\mathcal{P}_{\texttt{base}}$, we implement a few-shot prompt adding 5 examples ($\mathcal{P}_{\texttt{5shot}}$) and a chain-of-thought prompt simply asking the model to produce intermediate reasoning ($\mathcal{P}_{\texttt{cot}}$). We also ablate $\mathcal{P}_{\texttt{base}}$ by removing extended guidelines ($\mathcal{P}_{\texttt{noguide}}$) and not asking for explanations ($\mathcal{P}_{\texttt{noreason}}$). The full prompts can be found in Appendix˜D.
@@ 5 Results
We first investigate the effect of prompting techniques in Section˜5.1. Next, we evaluate the LLM annotations using automatic metrics (Section˜5.2) and manually analyze the errors in the model outputs (Section˜5.3).
@@ 5.1 Prompting Techniques
We perform preliminary experiments on the D2T-Eval task $\mathcal{D}_{\texttt{dev}}$ set using open models (Llama 3.3 and DeepSeek-R1) to study the differences between prompting techniques. The results are shown in Table˜2.
@@
Prompt $F_{1}$ $\gamma$ #a/o $F_{1}$ $\gamma$ #a/o
@@
Figure 2: Comparison of LLMs using $\mathcal{P}_{\texttt{base}}$ and human annotators. For $F_{1}$ score, the shadow bar denotes its soft variant. The dashed horizontal lines for D2T-Eval denote agreement between our human annotators. Detailed results are included in Tables˜11, 12, 13 and 14.
Including detailed guidelines seems beneficial: omitting the guidelines ($\mathcal{P}_{\texttt{noguide}}$) lowers the performance of both models. In contrast, not letting the model explain the annotation ($\mathcal{P}_{\texttt{noreason}}$) does not have a substantial effect. For Llama 3.3, the chain-of-thought (CoT) prompting ($\mathcal{P}_{\texttt{cot}}$) makes it produce fewer annotations per example than the base variant (0.8 vs. 2.4), leading to lower F1 and $\gamma$ scores. Llama 3.3 with $\mathcal{P}_{\texttt{cot}}$ tends to “overthink” the annotations, deciding not to annotate cases of errors against which it can find some arguments.
Few-shot prompting ($\mathcal{P}_{\texttt{5shot}}$) brings ambivalent results, increasing Llama 3.3 scores but doing the opposite for DeepSeek-R1. This observation is aligned with guo2025deepseek, who note that few-shot prompting degrades the performance of DeepSeek-R1. Given these considerations, we decided to use $\mathcal{P}_{\texttt{base}}$ for further experiments.
@@
Next, we compare LLM and human annotations using the metrics described in Section˜3.3. The overall results for all tasks are given in Figure˜2. We provide detailed results for individual tasks in the Appendix F.
Figure 3: Confusion matrix for D2T-Eval (Contradictory, Not checkable, Misleading, Incoherent, Repetitive, Other), averaged across models (see Table˜5 for category descriptions).
@@
DeepSeek-R1 generally outperforms the same-sized Llama 3.3. Its superiority is most pronounced on D2T-Eval ($F_{1}$-score of 0.23 vs. 0.18, $\gamma$ score of 0.19 vs. 0.11). The same observation applies to OpenAI models, where o3-mini outperforms GPT-4o. A notable exception to this trend is the non-reasoning Claude 3.7 Sonnet, which scores mostly on par with o3-mini and excels at MT-Eval.
Model Cost (\$/1k out) Time (s/out)
crowdworkers 500 129.1
Llama 3.3 - 21.6
DeepSeek-R1 - 227.5
Claude 3.7 Sonnet 10.5 9.0
o3-mini 3.6 21.8
Table 3: Estimate of costs and time requirements on D2T-Eval: crowdworkers on Prolific, open models (Llama 3.3, DeepSeek-R1), and proprietary models (Claude 3.7 Sonnet, o3-mini).
@@
For D2T-Eval, we compare model results with an average IAA on a subset of examples annotated by two human annotators. Here, o3-mini, Claude 3.7 and Gemini 2.0 mostly reach or surpass human agreement. For Propaganda, the upper bound of IAA is the result of dasanmartino2019finegrained, who report $\gamma=0.31$ for annotators before consolidation. This score is substantially higher than LLMs (the best LLM score being $\gamma=0.16$ for Claude 3.7 Sonnet). However, this task used expert annotators and has the largest number of categories. The latter property is reflected in the large difference between the soft and hard $F_{1}$ scores. [^5] We omit the comparison with human IAA in MT-Eval. While the WMT24 dataset for MT-Eval contains examples annotated with a pair of annotators, these examples take up only a small fraction and exhibit high variance between language pairs.
@@
Confusion matrices (see Figure˜3 and Appendix˜F) suggest that the models tend mainly to confuse related categories, which may be related to ambiguity or subjective understanding of category definitions. The models also use a less diverse distribution of categories than human annotators.
LLMs are more cost- and time-efficient than human annotators.
An important factor when comparing LLMs and human annotators is efficiency with respect to cost and time per output. For D2T-Eval, crowdsourced annotation for 1k outputs costs approximately \$500, while annotating the same amount of outputs with the high-performance o3-mini LLM costs \$3.60 (see Table˜3). In terms of time, the crowdworkers take 129.1 seconds per output on average, which is better than DeepSeek-R1 70B running on our local infrastructure, but an order of magnitude slower than the API-based models. [^6] Note that we do not ask the crowdworkers to give us a reason $r$ for the annotation, which would arguably make the responses of the crowdworkers slower. Therefore, LLMs are a more efficient alternative in terms of costs and time.
Figure 4: Results of our manual analysis. We analyzed 18 annotations and their explanations for each model and task (216 annotations in total). The color bars show annotations that we classified as Correct, Partially correct, Wrong category, Incorrect, and Undecidable. Detailed results are provided in Tables˜19 and 20.
@@ 5.3 Manual Analysis of LLM Annotations
To gain more insights into the qualitative aspects of LLM annotations, we manually analyzed the quality of LLM annotations on 216 samples from D2T-Eval and Propaganda. [^7] We do not include MT-Eval in the manual analysis due to our insufficient expertise in most target languages. For each model, we sampled three annotations per category in D2T-Eval and one annotation per category in Propaganda. Without access to the annotation source, we classified the annotations and their explanations as Correct, Partially correct, Wrong category, Incorrect, and Undecidable.
We show the results in Figure˜4 and Tables˜19 and 20. In total, we marked 49.5% of LLM-generated annotations and 50.5% of reasons as correct (with 9.2% of annotations and 12.5% of reasons additionally marked as partially correct). Reasoning models perform better, with 56.4% of their annotations and 58.3% reasons marked as completely correct. The most accurate annotations on D2T-Eval were those made by Gemini 2.0 and DeepSeek-R1. o3-mini performed well on both tasks, although Propaganda proved challenging for all models.
@@
We find that the models often select wrong error categories despite identifying real issues (e.g., labeling Contradictory statements as Incoherent). Models also tend to be overly attentive, flagging noise in the data (e.g., markup or off-topic content in Propaganda) as errors, or marking slight numerical variations (such as rounded values) as misleading. All of these cases could be tackled by more descriptive guidelines or additional examples. However, in some cases, the models also misread or misinterpreted the data (e.g., claiming wind speed measurements do not exist when they do), which hints at deeper issues with understanding the data. Incorrect explanations vary from incomplete explanations (addressing only part of a multi-issue span), irrelevant explanations (e.g., appealing to facts that are “missing” from the text) to incorrectly flagging subjective statements (e.g., “long-lasting usage”) as factual errors. Occasionally, the model admits that it marked a correct span as an error, such as in “The description of the game’s duration aligns with the data, providing coherent information”. [^8] This typically happened to GPT-4o, even though OpenAI API ensures JSON key ordering so the explanation should have been generated before the annotation (cf. Section 3.2).
@@
Concerningly, the LLM annotations that were marked as correct have only 24% hard character-level overlap (51% soft) with human annotations. This fact led us to analyze the quality of human annotations in D2T-Eval (the task in which we had the necessary domain expertise). Using the same methodology as we used for the LLM annotations, we annotated a limited sample of 108 human annotations. We marked 45.3% of the annotations as Correct, which is comparable to the LLM annotations (see Table˜4 for the results). These findings suggest that the task is hard even for human annotators, and the quality of annotations from crowdworkers varies, even if they are preselected using a qualification task.
@@ 6 Discussion
Here, we summarize our findings and discuss the implications of our results.
@@
The IAA between LLMs and human annotators is only moderate, suggesting LLMs cannot straightforwardly replace human annotators. However, using LLMs may be a reasonable option in scenarios such as D2T-Eval based on crowdworkers, where the strongest LLMs reach the average IAA between human annotators themselves. In other cases, when deciding whether to employ LLMs as span annotators, one needs to balance desired output quality with other practical aspects. Here, LLMs provide better flexibility, shorter response times, and lower costs. One should also consider the quality of available human annotators, as even qualified crowdworkers make similar amounts of errors as LLMs. It can be also assumed that LLM-based span annotation will benefit from future increases in LLM capabilities, while crowdworkers may increasingly rely on LLM to complete tasks veselovsky2023artificial. A promising solution seems to be a hybrid approach in which LLMs pre-annotate the text and humans post-edit the annotations zouhar2024aiassisted.
How to deploy LLMs as span annotators?
We recommend providing LLMs with detailed guidelines that describe conventions and how to handle ambiguous cases (cf. Figure˜5). In contrast, we do not recommend providing specific examples (cf. Figure˜8), as this approach did not bring consistent improvements. Arguably, this is due to the length and complexity of the examples, making them distracting to the model. When using LLMs with custom categories or guidelines, we recommend validating the model’s annotations against examples hand-annotated by experts on a sample of the data. In general, reasoning models tend to provide more reliable annotations at the cost of higher response times and token count.
Annotations
Source C P W I U
Human annotators 49 4 17 31 7
Table 4: Manual evaluation results for human annotators on D2T-Eval. Categories: C=Correct, P=Partially correct, W=Wrong category, I=Incorrect, U=Undecidable.
7 Conclusion
We showed that LLMs can serve as span annotators for three span annotation tasks: evaluating data-to-text generation, identifying errors in machine translation, and detecting propaganda in human-written texts. Our experiments show that LLMs achieve moderate agreement with skilled human annotators. The models perform best in D2T-Eval, where they are comparable to qualified crowdworkers. Reasoning models consistently outperform their instruction-tuned counterparts, delivering more accurate annotations and providing more valid explanations for their decisions. Automating span annotation with LLMs seems to be a promising alternative to fine-grained human evaluation sourced from crowdsourcing platforms, opening the way towards scalable and actionable automatic NLG evaluation methods.